OpenStudy (anonymous):

HELP Find the inverse of f(x) = 4x + 12

hartnn (hartnn):

let y= 4x+12 can you isolate 'x' from this equation ?

hartnn (hartnn):

start by subtracting 12 from both sides

OpenStudy (anonymous):

-8x ?

hartnn (hartnn):

nopes. 4x and 12 are not like terms, you cannot combine them subtracting 12 from both sides of y=4x+12 will just give you y-12 =4x got this ?

OpenStudy (anonymous):

ohh okay I understand that.

hartnn (hartnn):

now divide both sides by 4 what u get ?

OpenStudy (anonymous):

you get 1 and 3

hartnn (hartnn):

12/4 = 3 is correct, but when you divide y-12 by 4, you get \(\dfrac{y-12}{4}= \dfrac{y}{4}-\dfrac{12}{4}=\dfrac{y}{4}-3=\dfrac{4x}{4}=x\) got this ?

OpenStudy (anonymous):

yeahh I got it.

hartnn (hartnn):

last step is to just exchange x and y, and you get your inverse! \(f^{-1}(x)=\dfrac{x}{4}-3\) thats it ...ask if any doubt in any step.

OpenStudy (anonymous):

wait so whats x ?

OpenStudy (anonymous):

and f

hartnn (hartnn):

you need inverse of x, which we got as x/4 -3

OpenStudy (anonymous):

right, but how do you find the answer ? just change them ad solve

hartnn (hartnn):

go through the post again, all the steps are there, to get the final answer as x/4-3

OpenStudy (anonymous):

You said the last step is to exchange x and y so is it this ? x/-3-4

OpenStudy (anonymous):

I don't get that

hartnn (hartnn):

ok, we had x = y/4 - 3 exchanging x and y we get y= x/4 -3 which is your inverse function! :)

OpenStudy (anonymous):

ohh that's simple I was way thinking this

hartnn (hartnn):

yes, it is simple :) especially afte 2 or 3 question, you'll be able to do it mentally ;)

OpenStudy (anonymous):

haha I hope so

OpenStudy (anonymous):

I just tried that answer and it was wrong . /:

hartnn (hartnn):

x/4 -3 is not taken ? then lets not simplify it and keep it as (x-12) /4 put this.

hartnn (hartnn):

but in usual case, we need to simplify and so, x/4-3 should have been taken as correct :\

OpenStudy (anonymous):

y = x / 4 - 3 I put this and I only have one more attempt to get it right . so should I type it in google ?

hartnn (hartnn):

u sure the question is f(x) = 4x + 12 ? then i would say, try (x-12) /4

OpenStudy (anonymous):

Yes I'm positive!!

hartnn (hartnn):

then i am positive too, that the answer is (x-12)/4

OpenStudy (anonymous):

Ok i'll try it I don't put x = ( x - 12 ) / 4 ? It says Please put a single space between terms to increase the chance of auto grade working. Such as y = x + 3 instead of y=x+3.

hartnn (hartnn):

x = ( x - 12 ) / 4 ----- no inverse of f(x) = ( x - 12 ) / 4 or \(f^{-1}(x)\) = ( x - 12 ) / 4

OpenStudy (anonymous):

ok I will try it .

OpenStudy (anonymous):

do u know how to do this ? The amount of money collected by a toll booth is represented by the function A(m) = $2.50m where m stands for the number of minutes after midnight. What would be an appropriate domain for this function? Write your answer as an inequality and no spaces.

OpenStudy (anonymous):

I got it wrong again !!